Solve the system using elimination or substitution method.
1/4x +1/5y = 3/5
3x-6y=24

Answer by mananth(16946)

(1/4)x +(1/5)y = 3/5
LCD = 20
multiply by LCD
1*20/4x +1*20/5y = 3*20/5
5x+4y=12
3x-6y=24
5 x + 4 y = 12 .............1
-6 y = 24 -3 x ---------2
/ -6 =
y = -4 + 0.5 x
Substitute y in (1)
5 x + 4 ( -4 + 0.5 x
5 x -16 + 2 x = 12
5 x + 2 x = 12 16
7 x = 28
/ 7
x= 4
Plug the value of x in (1)
5 x + 4 y = 12
5 * 4 + 4 y = 12
20 + 4 y = 12
4 y = -8
/ 4
y= -2
m.ananth@hotmail.ca